OpenStudy (exceptionlearner):

Solve (x-2)4 (cube)=81

OpenStudy (aaronandyson):

take a screenshot of the original question.I am having trouble intercept the problem :/

OpenStudy (anonymous):

Can we assume that because you haven't answered that Directrix got it right? ^

OpenStudy (exceptionlearner):

Where's the draw tab?

OpenStudy (exceptionlearner):

Yes he got is some what but there's no 3 and he forgot "=81" at the end

OpenStudy (anonymous):

The draw tab will be under the text box that you are typing in,

OpenStudy (exceptionlearner):

Well I'm using a phone, I don't think they have draw tabs for it :(

OpenStudy (anonymous):

Well, if there is no 3, then what does the cube mean.

OpenStudy (exceptionlearner):

Sorry I got it mixed with another problem. I mean 4 to the fourth power

OpenStudy (exceptionlearner):

(x-2)4 (four to the fourth power)= 81

OpenStudy (anonymous):

Now we are getting somewhere.

OpenStudy (mathmale):

(x-2)4 (cube)=81 should be handwritten (use the Draw utility, below) , or typed in as (x-2)^4 = 81. If you want to be fancy, use Equation Editor, below:\[(x-2)^4=81\]

OpenStudy (mathmale):

If you take the fourth root of the left side of this equation, you will have isolated (x-2). You must also take the fourth root of 81, which, by the way, has more than one answer. Find all values of x for which \[(x-2)^4=81\]

OpenStudy (mathmale):

is true.

OpenStudy (exceptionlearner):

@math make that's exactly what the problem looks like. I'm new to openstudy

OpenStudy (exceptionlearner):

@mathmale **

OpenStudy (mathmale):

I'd be glad to help, but am unsure of what you're asking me for.

OpenStudy (exceptionlearner):

I need help solving the problem. That's all the problem ask Solve (x-2)^4=81

OpenStudy (mathmale):

"Solving" an equation implies the need to isolate the unknown. In this case, that unknown is x. What's the value (or, what are the values) of x that satisfy the original equation?

OpenStudy (mathmale):

Just as an example: let x=1 and test this to see whether it is indeed a solution of the given equation. Solve (x-2)4 (cube)=81 Let x=1, which leads to (1-2)^4=81. that's (-1)^4 = 81, or 1=81. Obviously not true, and so x=1 is not a solution.

OpenStudy (mathmale):

As before, this equation should be displayed as \[(x-2)^4=81.\]

OpenStudy (mathmale):

We need to isolate x-2, and then, in turn, to isolate x. Taking the 4th root of both sides of the equation:\[\sqrt[4]{(x-2)^4}=\pm \sqrt[4]{81}\]

OpenStudy (mathmale):

the left side of the equation becomes simply x-2. The right side becomes\[\pm \sqrt[4]{3^4}, \]

OpenStudy (mathmale):

since 3^4 = 81 = 3*3*3*3

OpenStudy (mathmale):

so we end up with \[x-2=\pm2\]

OpenStudy (mathmale):

Solve this for x. We end up with what? Note that because 81= 3^4, the maximum number of solutions could be 4. We're not finished until we've checked both x=3 and x=-3. Try it. Write the set that represents all of the solutions.

OpenStudy (mathmale):